There were 20 girls and 10 boys on the swim team last Summer. This summer,the number of girls decreased by 30% and the number of boys increased by 40%.How
many member are on the swim team this summer?

Answer by stanbon(75887)

There were 20 girls and 10 boys on the swim team last Summer. This summer,the number of girls decreased by 30% and the number of boys increased by 40%.How
many member are on the swim team this summer?
------
# = 0.7*20 + 1.4*10 = 14+14 = 28
-------------
Cheers,
Stan H.
-----------